What is the percent mass of calcium supplement if 1.37 g of calcium carbonate from a2.12g pill
Answer #1

2.12 g pill

CaCO3 = mass of Ca / total mass = 40 / (100) = 0.4

then

1.37 g = 1.37*0.4 = 0.548 g of Calcium

then

%mass of Ca = Ca/totla Mass * 100 = 0.548/(2.12) *100% = 25.84905 % of Ca in sample